OpenStudy (wolf):

y=x^2-2x-15 how do you find the vertex?

OpenStudy (wolf):

and the x intercepts are at 5 and -3?

OpenStudy (anonymous):

begin by completing the square

OpenStudy (anonymous):

to find the 'x' coordinate of the vertex, use x = -b/2a or similarly take the first derivative, set equal to zero and solve for 'x'. To find the 'y' coord. just plug the 'x' value into the original equation.

OpenStudy (anonymous):

Yes, the 'x' intercepts are 5 and -3: (x+3)(x-5) = 0

OpenStudy (wolf):

and vertex at 1,-16

OpenStudy (anonymous):

if you complete the square you'll arrive at \[y=(x-1)^{2}+13\] This is the standard form for a parabola. From this form, we then know that the vertex is at (1, 13)

OpenStudy (wolf):

oops i messed up haha

OpenStudy (anonymous):

oops, me too. Sorry

OpenStudy (wolf):

-1,-12?

OpenStudy (anonymous):

y=(x−1)2+14 so the vertex is at (1, 14)

OpenStudy (wolf):

obviously the derivative thing isnt working for me ahha

OpenStudy (anonymous):

that should work too. Take the first derivative: dy/dx=2x-2 set the first derivative equal to zero: 2x-2=0, x = 1

OpenStudy (anonymous):

plug in x= 1 to the original equation to find the y-coordinate of the vertex

OpenStudy (anonymous):

I got the coordinates of the vertex to be (1,-16)... After plugging in x=1. I may be wrong.

OpenStudy (wolf):

but when i plug 1 in i get -16?

OpenStudy (wolf):

which is what i had the first time before i got confused

OpenStudy (anonymous):

y(1)=\[y(1) = 1^{2}-2+15=14\]

OpenStudy (wolf):

-15

OpenStudy (wolf):

thats what you were doing wrong jjanelle haha

OpenStudy (anonymous):

1-2+15=14

OpenStudy (anonymous):

Oh, janelle, you got the -15 mixed up

OpenStudy (wolf):

hahah thank yall

OpenStudy (anonymous):

I can't read, good call.

OpenStudy (anonymous):

:)